OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form. Answer y = 7x + 18 y = 7x + 10 y = 7x – 18 y = 7x – 10

jimthompson5910 (jim_thompson5910):

slope m = (y2-y1)/(x2-x1) m = ( -3- 4)/(1-2) m = (-7)/(-1) m = 7 -------------- equation: y = mx+b 4 = (7)(2)+b 4 = 14+b 4-14 = b -10 = b b = -10 So the equation of the line is y = 7x-10 which is choice D

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw, glad to be of help